Carol brought a coloring book. It is 21 yd long and 17 yd wide. What is the diagonal length of the book?
Answer by Boreal(15235)

Pythagorean Theorem (drawing this helps)
side^2+side^2=hypotenuse^2
21^2+17^2=h^2
441+289=h^2=730.
h=27.02 yds diagonal
That's long for a book, so the units might be checked to see if they are really yards.